# Implicit Third-Order Peer Triplets with Variable Stepsizes for Gradient-Based Solutions in Large-Scale ODE-Constrained Optimal Control

### v1

@@ 7 A posteriori equi-distribution of global errors
where $\delta\in[0,1]$ is chosen as a weighting factor. Taylor expansions with exact solutions $(y^{\star}(t),p^{\star}(t))$ at off-step points $t=t_{ni}$ yield $\varepsilon^{Y}_{n}=h_{n}^{3}y^{\star(3)}(t_{n})+O(h_{n}^{4})$ and $\varepsilon^{P}_{n}=h_{n}^{3}p^{\star(3)}(t_{n})+O(h_{n}^{4})$. Supposing an asymptotic behaviour of the global errors,
$\|y(t_{n})-Y(t_{n})\|=O\left(h_{n}^{3}\|y^{\star(3)}(t_{n})\|\right),\;\|p(t_{n})-P(t_{n})\|=O\left(h_{n}^{3}\|p^{\star(3)}(t_{n})\|\right)$
with numerical approximations $Y(t_{n})=e_{1}^{\sf T}V_{4}^{-1}Y_{n}$ and $P(t_{n})=e_{1}^{\sf T}V_{4}^{-1}P_{n}$, the estimators defined above can be used to construct a new time mesh with nearly equally distributed global errors. In order to simplify equi-distribution of these errors, often linearized versions such as $h_{n}\|y^{\star(3)}(t_{n})\|^{1/3}$ and $h_{n}\|p^{\star(3)}(t_{n})\|^{1/3}$ are considered, see e.g. [14]. Defining the extended vector of approximate third derivatives $d_{n}:=((\varepsilon^{Y}_{n})^{\sf T},(\varepsilon^{P}_{n})^{\sf T})^{\sf T}/h_{n}^{3}$ and a piecewise constant mesh density function $\psi(t):=\|d_{n}\|^{1/3},\,t\in[t_{n},t_{n+1})$, on the old grid it holds that
@@
According to [14, Chapter 2.2.2] such a mesh function may be computed as the solution of the nonlinear boundary value problem
@@
For the solution of (60), we use MMPDELab by W. Huang [13] and start with a uniform grid. Linear finite elements with $N$ inner nodes $\xi_{n}$, $n=1,\ldots,N$, and a pseudo-timestepping scheme are applied. Additionally, we apply the smoothing of the mesh density function $\psi(t)$ in MMPDELab by controlling the ratios $\psi^{\prime}(t_{n}^{\prime})/\psi(t_{n}^{\prime})$ from (63) in order to guarantee $\sigma_{n}^{\prime}=1+\eta_{n}^{\prime}h_{n}^{\prime}$ with $|\eta_{n}^{\prime}|\leq 15$ on the new mesh $t_{n}^{\prime}=x(\xi_{n})$. Such meshes yield smoothly varying stepsize sequences and work quite well for our variable stepsize Peer methods.
@@ 8.1 Boundary control for the 1D Heat equation
where $\eta_{k}(0)=y^{\star}(0)^{\sf T}v^{[k]}$ and $\varphi_{1}(z):=(e^{z}-1)/z$. The exact control and adjoint are
@@
Figure 1: Dirichlet heat problem with $m=250$ spatial points. Examplary mesh density function $\psi(t)$ for AP4o33vgi, $N=63$ and adapted time grids for $N=15,31,63,123$ (top left). Convergence of the maximal control errors $\|U_{ni}-u(t_{ni})\|_{\infty}$ (top right), state errors $\|y(T)$-$y_{h}(T)\|_{\infty}$ (bottom left), and adjoint errors $\|p(0)$-$p_{h}(0)\|_{\infty}$ (bottom right) for uniform and adaptive time grids.
@@
The triangular approximation at the boundaries, i.e., using $\tilde{A}_{i}$ instead of $A_{i}$, $i=0,N$, requires 101510-15 iteration steps to meet a relative tolerance of $10^{-14}$ in the maximum norm for all smaller subsystems which are solved by the sparse direct solver of Matlab. For practical tolerances as $10^{-6}$, only 575-7 iteration steps are necessary, compared to one direct solve for the fully coupled (linear) system of fourfold size.
@@
where $m_{ref}>0$. $A$ and $\rho$ are constants that determine the rates of cell proliferation and cell death. We set $\rho=K_{\rho}/\bar{K}_{\rho}$ and $A=-K_{A}/\bar{K}_{A}$. The constants $\sigma_{l}$ and $\sigma_{r}$ are threshold and reference values for the nutrients, respectively. A characterization of all parameters and their values which are used in our computation are given in Appendix B.
@@
By using triangular approximations at the boundaries, it is still possible to apply a sparse direct solver to linear systems of size 196,609. Typical iteration counts are around 12 for a relative tolerance of $10^{-3}$, while approximately $6$ iterations are typically required for the inner steps.
@@ 9 Conclusion
We modified two of the third-order Peer two-step triplets from [19] and integrated automatic grid construction in efficient gradient-based solution algorithms for the fully coupled optimal control problem with unknown control $u(t)$. Positivity of the quadrature weights in all matrices $K$ is essential. To avoid additional memory requirements in large nonlinear systems, special triangular iterations with small contraction factors are applied in the initial and final steps. Both schemes, AP4o33vgi and AP4o33vsi, perform very well on a benchmark problem involving boundary heat control — an application where one-step methods typically suffer from order reduction, as demonstrated in [17]. Noteworthy, the symmetry properties of the pulcherrima scheme AP4o33vgi yield increased orders between three and four for all components $(y,u,p)$, even on coarse grids.
@@ Appendix A Coefficients of AP4o33vgi and AP4o33vsi
B: Model parameters for prostate cancer growth

### v2

@@ 7 A posteriori equi-distribution of global errors
where $\delta\in[0,1]$ is chosen as a weighting factor. In long-standing numerical experiments with Peer methods, we have discovered that the use of previous stage values, i.e., $\delta=0$, works quite reliable for stiff and very stiff problems. For mildly stiff problems, the choice $\delta=1$ often leads to a slightly better performance. For our examples in Section 8, we will present results for $\delta=0$.
Taylor expansions with exact solutions $(y^{\star}(t),p^{\star}(t))$ at off-step points $t=t_{ni}$ yield $\varepsilon^{Y}_{n}=h_{n}^{3}(y^{\star})^{(3)}(t_{n})+O(h_{n}^{4})$ and $\varepsilon^{P}_{n}=h_{n}^{3}(p^{\star})^{(3)}(t_{n})+O(h_{n}^{4})$. Supposing an asymptotic behaviour of the global errors,
$\|y(t_{n})-Y(t_{n})\|=O\left(h_{n}^{3}\|(y^{\star})^{(3)}(t_{n})\|\right),\;\|p(t_{n})-P(t_{n})\|=O\left(h_{n}^{3}\|(p^{\star})^{(3)}(t_{n})\|\right)$
with numerical approximations $Y(t_{n})=e_{1}^{\sf T}V_{4}^{-1}Y_{n}$ and $P(t_{n})=e_{1}^{\sf T}V_{4}^{-1}P_{n}$, the estimators defined above can be used to construct a new time mesh with nearly equally distributed global errors. In order to simplify equi-distribution of these errors, often linearized versions such as $h_{n}\|(y^{\star})^{(3)}(t_{n})\|^{1/3}$ and $h_{n}\|(p^{\star})^{(3)}(t_{n})\|^{1/3}$ are considered, see e.g. [13]. Defining the extended vector of approximate third derivatives $d_{n}:=((\varepsilon^{Y}_{n})^{\sf T},(\varepsilon^{P}_{n})^{\sf T})^{\sf T}/h_{n}^{3}$ and a piecewise constant mesh density function $\psi(t):=\|d_{n}\|^{1/3},\,t\in[t_{n},t_{n+1})$, on the old grid it holds that
@@
According to [13, Chapter 2.2.2] such a mesh function may be computed as the solution of the nonlinear boundary value problem
@@
For the solution of (71), we use MMPDELab by W. Huang [14] and start with a uniform grid. Linear finite elements with $N$ inner nodes $\xi_{n}$, $n=1,\ldots,N$, and a pseudo-timestepping scheme are applied. Additionally, we apply the smoothing of the mesh density function $\psi(t)$ in MMPDELab by controlling the ratios $\psi^{\prime}(t_{n}^{\prime})/\psi(t_{n}^{\prime})$ from (74) in order to guarantee $\sigma_{n}^{\prime}=1+\eta_{n}^{\prime}h_{n}^{\prime}$ with $|\eta_{n}^{\prime}|\leq 15$ on the new mesh $t_{n}^{\prime}=x(\xi_{n})$. Such meshes yield smoothly varying stepsize sequences and work quite well for our variable stepsize Peer methods. Good starting values for the new control vector $U^{\prime}$ at $\{t_{ni}^{\prime}\}$ can be computed using piecewise cubic polynomial interpolation of $U$. A pseudocode for the overall algorithm with the key parameters is given in Table 3.
@@
Algorithm: Adaptive Time Grids
1. Solve optimal control problem on uniform grid $\{t_{n}\}$ for control $U$.
2. Given weight $\delta\in[0,1]$, compute estimators $\varepsilon^{Y}_{i}$, $\varepsilon^{P}_{i}$, $i=0,\ldots,N$, of the global errors.
3. Given scaling parameters $atol_{Y}$, $atol_{P}$, $rtol_{Y}$, and $rtol_{P}$, build mesh density function $\psi(t)$.
4. Given smoothness parameter $\eta>0$, solve mesh PDE with smoothing of $\psi(t)$ such that $\sigma^{\prime}_{n}=1+\eta^{\prime}_{n}h^{\prime}_{n}$ with $|\eta^{\prime}_{n}|\leq\eta$ for the new grid $\{t^{\prime}_{n}\}$.
5. Update grid from $\{t_{n}\}$ to $\{t^{\prime}_{n}\}$.
6. Calculate initial guess for $U^{\prime}$ at $\{t_{ni}^{\prime}\}$ by interpolating $U$.
7. Solve optimal control problem on grid $\{t^{\prime}_{n}\}$.
Table 3: Algorithm for the adaptive time-grid construction by a posteriori equi-distribution of global errors.
@@ 8.1 Boundary control for the 1D Heat equation
where $\eta_{k}(0)=(y^{\star}(0))^{\sf T}v^{[k]}$ and $\varphi_{1}(z):=(e^{z}-1)/z$. The exact control and adjoint are
@@
Figure 1: Dirichlet heat problem with $m=250$ spatial points. Exemplary mesh density function $\psi(t)$ for AP4o33vgi, $N=63$ and adapted time grids for $N=15,31,63,123$ (top left). Convergence of the maximal control errors $\|U_{ni}-u(t_{ni})\|_{\infty}$ (top right), state errors $\|y(T)$-$y_{h}(T)\|_{\infty}$ (bottom left), and adjoint errors $\|p(0)$-$p_{h}(0)\|_{\infty}$ (bottom right) for uniform and adaptive time grids.
@@
The triangular approximation at the boundaries, i.e., using $\tilde{A}_{i}$ instead of $A_{i}$, $i=0,N$, requires 101510-15 iteration steps of the form (45) to meet a relative tolerance of $10^{-14}$ in the maximum norm for all four smaller subsystems which are solved by the sparse direct solver of Matlab. For practical tolerances as $10^{-6}$, only 575-7 iteration steps are necessary. This results in approximately $25$ linear direct solves of size $m=250$ for each boundary step instead of one direct solve for the fully coupled (linear) systems written out as $(A_{i}\otimes I_{m}-h_{i}K\otimes{\bf J}_{i})Y_{i}=b_{i}$ and $(A_{i}^{\sf T}\otimes I_{m}-h_{i}K\otimes{\bf J}_{i}^{\sf T})P_{i}={b}_{i}^{\dagger}$, $i=0,N$, of size 4m=1,0004*m=1,000 with full matrices $A_{i}\in{\mathbb{R}}^{4\times 4}$.
Remark 8.1
It is a known fact that one-step methods such as Rosenbrock, Runge–Kutta, and extrapolation methods suffer from order reduction when applied to stiff ODEs or semi-discrete nonlinear parabolic PDEs [22, 23]. In the latter case, the temporal order of convergence is mainly influenced by the spatial regularity of the solution, which usually depends on the boundary conditions. Time-dependent boundary conditions constitute the most severe case. Thus, classical symplectic integrators, such as the A-stable Runge-Kutta-Gauss methods, lose their advantage of high-order accuracy when applied to boundary control problems. For example, the two-stage fourth-order Runge-Kutta-Gauss method shows a reduction to first order for the approximation of the control. The interested reader is referred to [17] for a detailed discussion of this phenomenon.
@@
where $m_{ref}>0$. $A$ and $\rho$ are constants that determine the rates of cell proliferation and cell death. We set $\rho=K_{\rho}/\bar{K}_{\rho}$ and $A=-K_{A}/\bar{K}_{A}$. The constants $\sigma_{l}$ and $\sigma_{r}$ are threshold and reference values for the nutrients, respectively. A characterization of all parameters and their values which are used in our computation are given in Appendix C.
@@
By using triangular approximations at the boundaries, i.e., replacing $A_{i}$ by $\tilde{A}_{i}$, $i=0,N$, and solving linear systems of the form (45), it is still possible to apply a sparse direct solver to linear systems of size $196,609$. Typical iteration counts are around 12 for a relative tolerance of $10^{-3}$, while approximately $6$ Newton iterations are typically required for the four nonlinear subsystems of the standard Peer steps with $i=1,\ldots,N-1$. This means that the strong coupling of the stage solutions at the boundaries resulting in nonlinear systems of size $786,436$ can be resolved by solving a sequence of smaller systems of size $196,609$ at the expense of only doubling the computational effort in the boundary steps.
@@ 9 Conclusion
We modified two of the third-order Peer two-step triplets from [19] and integrated automatic grid construction in efficient gradient-based solution algorithms for the fully coupled optimal control problem with unknown control $u(t)$. Positivity of the quadrature weights in all matrices $K$ is essential. To avoid additional memory requirements in large nonlinear systems, special triangular iterations with small contraction factors are applied in the initial and final steps. The pulcherrima triplet AP4o33vgi possesses LSRK-form (36) for state and adjoint equations, yielding super-convergence order three for arbitrary stepsize ratios. An additional grid-smoothness condition $|\sigma_{n}-1|\leq\eta h_{n}$ is necessary to ensure third-order adjoint super-convergence for the Peer triplet AP4o33vsi. Extensive numerical experiments showed that $\eta\leq 15$ is an appropriate choice for the design of smooth adaptive time grids. This moderate restriction comes with an significantly increased stability angle of $83.74^{o}$, compared to $61.59^{o}$ of the pulcherrima triplet.
Both schemes, AP4o33vgi and AP4o33vsi, perform very well on a benchmark problem involving boundary heat control — an application where one-step methods typically suffer from order reduction, as demonstrated in [17]. Noteworthy, the symmetry properties of the pulcherrima scheme AP4o33vgi yield increased orders between three and four for all components $(y,u,p)$, even on coarse grids.
@@
In both applications, the post-processed control $U^{\ddagger}$ introduced in (17) offers no discernible benefit. We infer that for Peer methods, which perform close to their theoretical order or even better, the approximation quality of $U$ is nearly optimal and postprocessing is not advisable in general.
@@ Appendix A Coefficients of AP4o33vgi and AP4o33vsi
Appendix B List of frequently used symbols
$A,B(\sigma),K$ coefficients of the standard method, $A$ triangular, $K\succ 0$ diagonal
$A_{0},a,A_{N},w$, $\tilde{A}_{0},\tilde{A}_{N}$ coefficients of the two boundary methods and triangular approximations for iterative solution
$\bar{B}_{n}=A_{n}^{-1}B_{n},\,\tilde{B}_{n+1}^{\sf T}=A_{n}^{\sf-T}B_{n+1}^{\sf T}$ corresponding zero-stability matrices
$Y_{n}=\big(Y_{ni}\big)_{i=1}^{s}$, $U_{n},P_{n}$ stacked numerical stages of state, control, and Lagrange multipliers
$t_{n},h_{n},\sigma_{n}\in[\underline{\sigma},\bar{\sigma}]$ gridpoints, stepsizes, and stepsize ratios $\sigma_{n}=h_{n}/h_{n-1}$
${\bf c},V_{q},{\cal P}_{q}=\exp(\tilde{E}_{q})$ node vector, $s\times q$-Vandermonde matrix, $q\times q$-Pascal matrix
$S_{n,q}=\,\mbox{diag}_{i=1}^{q}(\sigma_{n}^{i-1})$ stretching matrix for Taylor coefficients
${\cal Q}_{q,q}=V_{q}^{\sf T}B(\sigma)V_{q}{\cal P}_{q}^{-1}$ plays central role, order conditions imply ${\cal Q}_{q,q}=e_{1}e_{1}^{\sf T}$, $q<s$
$\hat{A}_{n}=V_{s}^{\sf T}A_{n}V_{s}$, $\hat{K},\hat{B}(\sigma)$ congruent matrices having restricted form related to ${\cal Q}_{q,q}$
$W\in{\mathbb{R}}^{s\times s}$ fixed weight matrix which block-diagonalizes all $\bar{B}(\sigma)=A^{-1}B(\sigma)$
$\tau_{n}^{Y},\tau_{n}^{P}$ local errors of state and adjoint variables, leading terms $\beta_{q,n}(\sigma),\beta_{q,n}^{\dagger}(\sigma)$
$err_{q,n}$, $err_{q,n}^{\dagger}$ leading error constants for Peer method ($\|\beta_{q,n}(1)\|_{\infty}$), and adjoint
$\|\phi^{(q:r)}\|_{[n]}$ local supremum norm on 3 adjacent subintervals of $\phi^{(k)}$, $q\leq k\leq r$
$\varepsilon_{n}^{Y},\varepsilon_{n}^{P}$, $v_{1}^{\sf T},v_{2}^{\sf T}$ a-posteriori error estimates and its coefficients
$\psi(x)$ mesh density function for grid optimization
Appendix C Model parameters for prostate cancer growth